An investor has GH¢40M to invest for a year at a simple interest rate. If part is invested at 8% and the rest at 12%, how much should be invested at each rate to yield 11% on the amount invested?
Answer #1

Let x millions has been invested in 8% rate. Then:

0.08*x + 0.12*(40 - x) = 0.11*40

4.8 - 0.04x = 4.4

0.04x = 0.4

x = 10

So,

Investment in 8% rate = $10 million

Investment in 12% rate = $40 - $10 = $30 million